# Balanced rectangles over Sturmian words and minimal discrepancy intervals

### v1

@@ 6.3.  Technical lemmas for the case $x^{*}<n$
Assume $2\leq m\leq n$ and $q_{M-1}<m\leq q_{M}$ with $m\geq q_{2}$ (or $m\geq q_{1}$ if $\alpha<1/2$). Moreover, assume that $x^{*}<n$. Then $f$ is bijective in exactly the two following cases:
  • $\langle q_{M}\alpha\rangle<1/2$ and $m=q_{M}$;
  • $\langle q_{M}\alpha\rangle<1/2$, $x^{*}\geq n-q_{M-1}$, and $m=q_{M-2}+aq_{M-1}$ for some $1\leq a\leq a_{M}-1$ and $M-2\geq 0$ (or $M-2\geq 1$ if $\alpha>1/2$).
@@
We distinguish between five cases, based on whether $\langle q_{M}\alpha\rangle<1/2$ or $\langle q_{M-1}\alpha\rangle<1/2$, and some extra conditions.
Case 1: $\langle q_{M-1}\alpha\rangle<1/2$.
Then, by Lemma 6.9, we have $x(\ell)=x^{*}$ for $\ell\in R_{0}$ and $x(\ell)=x^{*}+q_{M-1}$ for $\ell\in R_{1}$. Since $m-1\geq q_{M-1}$ and $m+x^{*}-n-1+q_{M-1}\geq q_{M-1}$, the range $R_{0}\cup R_{1}$ contains at least $q_{M-1}+1$ consecutive integers. Therefore, there exist $\ell\in R_{0}$ and $\ell^{\prime}\in R_{1}$ such that $\ell^{\prime}-\ell=q_{M-1}$. For these $\ell,\ell^{\prime}$ we have $x(\ell^{\prime})-x(\ell)=q_{M-1}=\ell^{\prime}$-$\ell$, and so $f$ is not bijective by Lemma 6.3.
Case 2: $\langle q_{M}\alpha\rangle<1/2$ and $m=q_{M}$.
By Lemma 6.9, we have either $x(\ell)=x^{*}$ or $x(\ell)=x^{*}+q_{M}$ for all $k\in R_{0}\cup R_{1}^{\prime}=[0,m-1]=[0,q_{M}-1]$. Therefore, it is clear that we cannot have $\ell-\ell^{\prime}=x(\ell)-x(\ell^{\prime})$ for $\ell\neq\ell^{\prime}\in[0,m-1]$, and so $f$ is bijective by Lemma 6.3.
Case 3: $\langle q_{M}\alpha\rangle<1/2$ and $m=q_{M-2}+aq_{M-1}$ with $1\leq a\leq a_{M}-1$ and $x^{*}\geq n-q_{M-1}$.
Then, by Lemma 6.9, we have $x(\ell)=x^{*}$ for $\ell\in R_{0}$ and $x(\ell)=x^{*}+q_{M-2}+aq_{M-1}$ for $\ell\in R_{1}^{\prime\prime}$. Moreover, since we are assuming $x^{*}\geq n-q_{M-1}$, we have
$x^{*}-n+q_{M-2}+(a+1)q_{M-1}-1\geq q_{M-2}+aq_{M-1}-1=m-1.$
In other words, $R_{0}\cup R_{1}^{\prime\prime}=[0,m+1]$, and so for all $\ell\in[0,m-1]=[0,q_{M-2}+aq_{M-1}-1]$ either $x(\ell)=x^{*}$ or $x(\ell)=x^{*}+q_{M-2}+aq_{M-1}$. As in Case 2, it is clear that $f$ is bijective.
Case 4: $\langle q_{M}\alpha\rangle<1/2$ and $m=q_{M-2}+aq_{M-1}$ with $1\leq a\leq a_{M}-1$, but now $x^{*}<n-q_{M-1}$.
Then
$x^{*}-n+q_{M-2}+(a+1)q_{M-1}-1<m-1,$
and so there is at least one $\ell\in R_{2}^{\prime\prime}$. Now since the range $R_{1}^{\prime\prime}$ contains exactly $q_{M-1}$ integers (note that $m=q_{M-2}+aq_{M-1}$), the range $R_{1}^{\prime\prime}\cup R_{2}^{\prime\prime}$ contains at least $q_{M-1}+1$ consecutive integers. Moreover, Lemma 6.9 says that $x(\ell)=x^{*}+q_{M-2}+aq_{M-1}$ for $\ell\in R_{1}^{\prime\prime}$ and $x(\ell)=x^{*}+q_{M-2}+(a+1)q_{M-1}$ for $\ell\in R_{2}^{\prime\prime}$. Thus, there must exist $\ell\in R_{1}^{\prime\prime}$ and $\ell^{\prime}\in R_{2}^{\prime\prime}$ with $\ell^{\prime}-\ell=q_{M-1}=x(\ell^{\prime})-x(\ell)$, and so $f$ is not bijective.
Case 5: $\langle q_{M}\alpha\rangle<1/2$ and $q_{M-2}+aq_{M-1}<m<q_{M-2}+(a+1)q_{M-1}$ for some $0\leq a\leq a_{M}-1$. If the range $R_{2}^{\prime\prime}$ is non-empty, then $f$ is not bijective by the same argument as in Case 4. Otherwise, $R_{0}\cup R_{1}^{\prime\prime}=[0,m-1]$, and the range contains $m>q_{M-2}+aq_{M-1}$ consecutive integers. Thus, we can find $\ell\in R_{0}$, $\ell^{\prime}\in R_{1}^{\prime\prime}$ with $\ell^{\prime}-\ell=q_{M-2}+aq_{M-1}=x(\ell^{\prime})-x(\ell)$, and so $f$ is not bijective either. ∎
6.4.  The cases where $x^{*}\neq n$
@@
Let $2\leq m\leq n$. Assume that $q_{M-1}<m<q_{M}$ and that $m$ is not a semi-convergent. Moreover, assume that $n\neq x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}$. Then $f_{\textnormal{left}},f_{\textnormal{right}}$ are not bijective.
@@
Now, if $x_{\textnormal{left}}^{*}<n$, then Lemmas 6.8 and 6.10 imply that $f_{\textnormal{left}}$ is not bijective.
If $x_{\textnormal{left}}^{*}>n$, then by Lemma 6.7 we have $x_{\textnormal{right}}^{*}(m,q_{T}-n)<q_{T}-n$. Then again Lemmas 6.8 and 6.10 imply that the intervals of length $\left\{(q_{T}-n)\alpha\right\}$ are not balanced with respect to $(\alpha,m)$, and so neither are those of length $\left\{n\alpha\right\}$. ∎
@@
Choose $\langle\cdot\rangle\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ so that $\langle n^{[\geq M]}\rangle<1/2$. Note that $M\geq 2$ if $\alpha>1/2$, and $M\geq 1$ if $\alpha<1/2$. We want to show that $x^{*}<n$ because then we can finish with Lemma 6.10. In fact, we want to show that $x^{*}\leq n^{[\geq M]}$.
First, note that $n^{[\geq M]}=n-n^{[\leq M-1]}\in[n-m+1,n-1]$ because $0<n^{[\leq M-1]}<q_{M}=m$. Therefore, we want to show that for all $x\in[n^{[\geq M]}+1,n+m-1]$ we have $\langle x\alpha\rangle>\langle n^{[\geq M]}\alpha\rangle$.
If $k_{0}(n^{[\geq M]})\geq M+1$, then Lemma 5.13 tells us that $n^{[\geq M]}$ minimises $\langle x\alpha\rangle$ in the range $n^{[\geq M]},n^{[\geq M]}+q_{M}+q_{M+1}-1]$. The upper bound of this range is
@@
Therefore, we indeed have $x^{*}\leq n^{[\geq M]}$.
Now assume $k_{0}(n^{[\geq M]})=M$. By the same argument as before, Lemma 5.13 tells us that $n^{[\geq M]}$ minimises $\langle x\alpha\rangle$ in the range $n^{[\geq M]},n^{[\geq M]}+q_{M-1}+q_{M}-1]$. If
@@
then, as before, we know that $n^{[\geq M]}$ minimises $\langle x\alpha\rangle$ in the range $n^{[\geq M]},n+m-1]$, and thus $x^{*}\leq n^{[\geq M]}$. Assume that the inequality (6.4) does not hold, i.e., after cancelling $q_{M}-1=m-1$, we get
@@
Since $n^{[\leq M-1]}=n-n^{[\geq M]}$, this implies $n^{[\leq M-1]}>q_{M-1}$, and so by the property (5.6) of the Ostrowski representation, we have $b_{M-1}(n)>0$. But then, by the digit rules for Ostrowski representations, we must have $b_{M}(n)\leq a_{M+1}-1$, and so we can use the stronger statement in Lemma 5.13: In this case $n^{[\geq M]}$ minimises $\langle x\alpha\rangle$ in the range $n^{[\geq M]},n^{[\geq M]}+q_{M-1}+2q_{M}-1]$. This now covers the full range $n^{[\geq M]}+1,n+m-1]$, and thus $x^{*}\leq n^{[\geq M]}$.
Overall, we have proven that $x^{*}\leq n^{[\geq M]}<n$, and so Lemma 6.10 says that $f$ is bijective if and only if $\langle q_{M}\alpha\rangle<1/2$, which is equivalent to $k_{\geq M}(n)\equiv M\pmod{2}$ by Lemma 5.5. ∎
@@
Let $m=q_{M-2}+aq_{M-1}$ with $1\leq a\leq a_{M}-1$ and $M-2\geq 1$ (or $M-2\geq 0$ if $\alpha<1/2$). Let $n\geq m$ and assume that $n$ is not of the shape (a) or (b) from Lemma 6.5 (or, in other words, $x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}\neq n$). Then $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective if and only if $k_{\geq M-1}(n)\equiv M\pmod{2}$.
@@
Let $m,n$ be as in the lemma. We distinguish between two cases according to whether $q_{M-1}$ shows up in the representation of $n$.
@@
Since $n^{[\leq M-2]}<q_{M-1}<m$, we have $n^{[\geq M]}\in[n-m+1,n+m-1]$. Moreover, since $m\leq q_{M}-q_{M-1}$ and $n^{[\leq M-2]}>n-q_{M-1}$, we have $n-m+1,n+m-1]\subseteq[n^{[\leq M-2]}-q_{M}+1,n^{[\leq M-2]}+q_{M}-1]$. Thus, Lemma 5.12 implies that $n^{[\geq M]}=x^{*}$ for the appropriate choice of $x^{*}\in\{x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}\}$. In particular, we have $x^{*}<n$. Now Lemma 6.10 says that $f$ is bijective if and only if $\langle q_{M}\alpha\rangle<1/2$, which is equivalent to $k_{\geq M-1}(n)=k_{0}(n^{[\geq M]})\equiv M\pmod{2}$, by Lemma 5.5.
@@
Our first goal is to show that we may assume that one of $x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}$ is smaller than $n$. If $x_{\textnormal{left}}^{*}<n$, this is of course the case. Assume now that $x_{\textnormal{left}}^{*}>n$. Then Lemma 6.7 says that if we consider $q_{T}-n$ instead of $n$ for sufficiently large $T$, we get $x_{\textnormal{right}}^{*}(m,q_{T}-n)<q_{T}-n$. Lemma 6.6 guarantees that the balancedness property doesn’t change if we replace $n$ by $q_{T}-n$. Lemma 5.18 implies (note that by assumption $n$ is not of the shape (b)) that $k_{\geq M-1}(q_{T}-n)\equiv M-1\pmod{2}$. If $q_{M-1}$ does not show up in the representation of $q_{T}-n$, then we know from Case 1 that we are in the not balanced situation. If $q_{M-1}$ shows up in the representation of $q_{T}-n$, then $q_{T}-n$ has the same shape (the shape of Case 2) as $n$, and we can replace $q_{T}-n$ by $n$, knowing that now the new $x_{\textnormal{right}}^{*}$ is smaller than the new $n$.
@@
Overall, in Case 2, we may now assume that at least one of $x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}$ is smaller than $n$. We fix this $x^{*}<n$, and our goal is to show that $f$ is not bijective.
@@
In order to apply Lemma 6.10 we need to use the fact that $\langle q_{M}\alpha\rangle<1/2$ if and only if $k_{0}(x^{*}\alpha)\equiv M\pmod{2}$. This is guaranteed by Lemma 5.5, unless $k_{0}(x^{*})=0$ (or $k_{0}(x^{*})=1$ if $\alpha<1/2$). In the exceptional cases we still have $\langle q_{0}\alpha\rangle<1/2$ if $k_{0}(x^{*})=0$ (and $\langle q_{1}\alpha\rangle<1/2$ if $\alpha<1/2$ and $k_{0}(x^{*})=1$). (To see this, assume the contrary and use Lemma 5.3 and the assumption $x^{*}<n$ to obtain a contradiction.)
@@
Now, if $k_{0}(x^{*})\equiv M+1\pmod{2}$, then Lemma 6.10 says that $f$ is not bijective, as desired.
@@
If $k_{0}(x^{*})\equiv M\pmod{2}$ and $x^{*}<n-q_{M-1}$, then Lemma 6.10 says the $f$ is not bijective as well.
@@
We are left with the case $n-q_{M-1}\leq x^{*}<n$ and $k_{0}(x^{*})\equiv M\pmod{2}$. We need to show that this is impossible for $n=n^{[\leq M-2]}+b_{M-1}q_{M-1}+n^{[\geq M]}$ with $1\leq b_{M-1}\leq a_{M}$ (and $n$ not of the shape (b)). The shape of $n$ and the assumptions $n-q_{M-1}\leq x^{*}<n$ and $k_{0}(x^{*})\equiv M\pmod{2}$ imply that $k_{0}(x^{*})\geq M$ is impossible. Thus, we must have $k_{0}(x^{*})\leq M-2$. But then, $\langle(x^{*}+q_{M-1})\alpha\rangle<\langle x^{*}\alpha\rangle$ by Lemma 5.9. Since $x^{*}+q_{M-1}\in[n-m+1,n+m-1]$, this contradicts the fact that $x^{*}$ minimises $\langle x\alpha\rangle$ in the range [nm+1,n+m1][n-m+1,n+m-1]. ∎
6.5.  Finishing the proof of the full characterisation

### v2

@@ 6.4.  The cases where $\widehat{x}_{*}\neq n$
Let $2\leq q_{1}\leq m\leq n$ and $q_{M-1}<m<q_{M}$. Assume that $\langle q_{M}\alpha\rangle_{*}<1/2$ and $\widehat{x}_{*}=n-q_{M-1}$. Then $n\in\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$.
@@
The assumptions $q_{1}\leq m<q_{M}$ imply $M\geq 2$. The assumption $\langle q_{M}\alpha\rangle_{*}<1/2$ implies that $k_{0}(\widehat{x}_{*})\equiv M\pmod{2}$ (or possibly $k_{0}(\widehat{x}_{*})=0$).
If $k_{0}(\widehat{x}_{*})\geq M$, then it follows from the basic properties of Ostrowski representations that $n=q_{M-1}+\widehat{x}_{*}$ has one of the two shapes (a), (b) from Lemma 6.5, which implies $n\in\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$.
If $k_{0}(\widehat{x}_{*})\leq M-2$, then Lemma 5.9 implies that $\langle n\alpha\rangle_{*}=\langle(\widehat{x}_{*}+q_{M-1})\alpha\rangle_{*}<\langle\widehat{x}_{*}\alpha\rangle_{*}$, contradicting the definition of $\widehat{x}_{*}$. ∎
Remark 6.11.
In view of Lemma 6.10, if we assume $n\notin\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$, then Lemma 6.9 now only gives us two cases where $f_{*}$ is bijective, namely where $m$ is either a convergent or a semi-convergent (and some extra condition).
@@
Let $2\leq q_{1}\leq m\leq n$. Assume that $q_{M-1}<m<q_{M}$ and that $m$ is not a semi-convergent. Moreover, assume that $n\notin\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$. Then $f_{\textnormal{left}},f_{\textnormal{right}}$ are not bijective.
@@
Now, if $\widehat{x}_{\textnormal{left}}<n$, then Lemmas 6.9 and 6.10 imply that $f_{\textnormal{left}}$ is not bijective.
If $\widehat{x}_{\textnormal{left}}>n$, then by Lemma 6.7 we have $\widehat{x}_{\textnormal{right}}(m,q_{T}-n)<q_{T}-n$. Then again Lemmas 6.9 and 6.10 imply that the intervals of length $\left\{(q_{T}-n)\alpha\right\}$ are not balanced with respect to $(\alpha,m)$, and so neither are those of length $\left\{n\alpha\right\}$. ∎
@@
Note that $M\geq 1$. Choose $\langle\cdot\rangle_{*}\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ so that $\langle n^{[\geq M]}\alpha\rangle_{*}<1/2$. We want to show that $\widehat{x}_{*}<n$ because then we can finish with Lemma 6.9. In fact, we want to show that $\widehat{x}_{*}\leq n^{[\geq M]}$.
First, note that $n^{[\geq M]}=n-n^{[\leq M-1]}\in[n-m+1,n-1]$ because $0<n^{[\leq M-1]}<q_{M}=m$. Therefore, we want to show that for all $x\in[n^{[\geq M]}+1,n+m-1]$ we have $\langle x\alpha\rangle_{*}>\langle n^{[\geq M]}\alpha\rangle_{*}$.
Assume first that $k_{0}(n^{[\geq M]})\geq M+1$. Then Lemma 5.13 tells us that $n^{[\geq M]}$ minimises $\langle x\alpha\rangle_{*}$ in the range $n^{[\geq M]},n^{[\geq M]}+q_{M}+q_{M+1}-1]$. The upper bound of this range is
@@
Therefore, we indeed have $\widehat{x}_{*}\leq n^{[\geq M]}$.
Now assume $k_{0}(n^{[\geq M]})=M$. By the same argument as before, Lemma 5.13 tells us that $n^{[\geq M]}$ minimises $\langle x\alpha\rangle_{*}$ in the range $n^{[\geq M]},n^{[\geq M]}+q_{M-1}+q_{M}-1]$. If
@@
then, as before, we know that $n^{[\geq M]}$ minimises $\langle x\alpha\rangle_{*}$ in the range $n^{[\geq M]},n+m-1]$, and thus $\widehat{x}_{*}\leq n^{[\geq M]}$. If the inequality (6.4) does not hold, then, after cancelling $q_{M}-1=m-1$, we get
@@
Since $n^{[\leq M-1]}=n-n^{[\geq M]}$, this implies $n^{[\leq M-1]}>q_{M-1}$, and so by the property (5.6) of Ostrowski representations, we have $b_{M-1}(n)>0$. But then, by the digit rules for Ostrowski representations, we must have $b_{M}(n)\leq a_{M+1}-1$, and so we can use the stronger statement in Lemma 5.13: In this case $n^{[\geq M]}$ minimises $\langle x\alpha\rangle_{*}$ in the range $n^{[\geq M]},n^{[\geq M]}+q_{M-1}+2q_{M}-1]$. This now covers the full range $n^{[\geq M]}+1,n+m-1]$, and thus $\widehat{x}_{*}\leq n^{[\geq M]}$.
Overall, we have proven that $\widehat{x}_{*}\leq n^{[\geq M]}<n$, and so Lemma 6.9 says that $f_{*}$ is bijective if and only if $\langle q_{M}\alpha\rangle_{*}<1/2$, which is equivalent to $k_{\geq M}(n)\equiv M\pmod{2}$ by Lemma 5.5. ∎
@@
Let $m=q_{M-2}+aq_{M-1}$ with $1\leq a\leq a_{M}-1$ and $M-2\geq 0$. Let $n\geq m$ and assume that $n$ is not of the shape (a) or (b) from Lemma 6.5 (or, in other words, $n\notin\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$). Then $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective if and only if $k_{\geq M-1}(n)\equiv M\pmod{2}$.
@@
Let $m,n$ be as in the lemma. We distinguish between two cases according to whether $q_{M-1}$ shows up in the representation of $n$ or not.
@@
Since $n^{[\leq M-2]}<q_{M-1}<m$, we have $n^{[\geq M]}\in[n-m+1,n+m-1]$.
[…]
@@
Our first goal is to show that we may assume that one of $\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}$ is smaller than $n$. If $\widehat{x}_{\textnormal{left}}<n$, this is of course the case. Assume now that $\widehat{x}_{\textnormal{left}}>n$. Then Lemma 6.7 says that if we consider $q_{T}-n$ instead of $n$ for sufficiently large $T$, we get $\widehat{x}_{\textnormal{right}}(m,q_{T}-n)<q_{T}-n$. Lemma 6.6 guarantees that the balancedness property doesn’t change if we replace $n$ by $q_{T}-n$. Lemma 5.17 implies (note that by assumption $n$ is not of the shape (b)) that $k_{\geq M-1}(q_{T}-n)\equiv M-1\pmod{2}$. If $q_{M-1}$ does not show up in the representation of $q_{T}-n$, then we know from Case 1 that we are in the not balanced situation. If $q_{M-1}$ shows up in the representation of $q_{T}-n$, then $q_{T}-n$ has the same shape (the shape of Case 2) as $n$, and we can replace $q_{T}-n$ by $n$, knowing that now the new $\widehat{x}_{\textnormal{right}}$ is smaller than the new $n$.
@@
Overall, in Case 2, we may now assume that at least one of $\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}$ is smaller than $n$. We fix this $\widehat{x}_{*}<n$, and our goal is to show that $f_{*}$ is not bijective.
@@
In order to apply Lemma 6.9 we use the fact that $\langle q_{M}\alpha\rangle_{*}<1/2$ if and only if $k_{0}(\widehat{x}_{*})\equiv M\pmod{2}$. This is guaranteed by Lemma 5.5, unless $k_{0}(\widehat{x}_{*})=0$. In the exceptional case we still have $\langle q_{0}\alpha\rangle_{*}<1/2$. (To see this, assume the contrary and use Lemma 5.3 and the assumption $\widehat{x}_{*}<n$ to obtain a contradiction.)
@@
Now, if $k_{0}(\widehat{x}_{*})\equiv M+1\pmod{2}$, i.e., $\langle q_{M-1}\alpha\rangle_{*}<1/2$, then Lemma 6.9 says that $f_{*}$ is not bijective, as desired.
@@
If $k_{0}(\widehat{x}_{*})\equiv M\pmod{2}$, i.e., $\langle q_{M}\alpha\rangle_{*}<1/2$, and if $\widehat{x}_{*}<n-q_{M-1}$, then Lemma 6.9 again says that $f_{*}$ is not bijective.
@@
We are left with the case $n-q_{M-1}\leq\widehat{x}_{*}<n$ and $k_{0}(\widehat{x}_{*})\equiv M\pmod{2}$. We need to show that this is impossible for $n=n^{[\leq M-2]}+b_{M-1}q_{M-1}+n^{[\geq M]}$ with $1\leq b_{M-1}\leq a_{M}$ (and $n$ not of the shape (b)). The shape of $n$ and the assumptions $n-q_{M-1}\leq\widehat{x}_{*}<n$ and $k_{0}(\widehat{x}_{*})\equiv M\pmod{2}$ imply that $k_{0}(\widehat{x}_{*})\geq M$ is impossible. Thus, we must have $k_{0}(\widehat{x}_{*})\leq M-2$. But then, $\langle(\widehat{x}_{*}+q_{M-1})\alpha\rangle_{*}<\langle\widehat{x}_{*}\alpha\rangle_{*}$ by Lemma 5.9. Since $\widehat{x}_{*}+q_{M-1}\in[n-m+1,n+m-1]$, this contradicts the fact that $\widehat{x}_{*}$ minimises $\langle x\alpha\rangle_{*}$ in the range [nm+1,n+m1][n-m+1,n+m-1]. ∎
6.5.  Very small $m$
In the previous two subsections we assumed $m\geq q_{1}$ for technical reasons. The cases where $2\leq m<q_{1}$ are actually quite easy and we deal with them by going back to the original definition of balanced intervals.
Lemma 6.15.
Let $2\leq m\leq n$ with $m<q_{1}$. Then the intervals of length $\left\{n\alpha\right\}$ are balanced with respect to $\{0,\alpha,\ldots,(m-1)\alpha\}$ if and only if either $n$ is either of the shape $n=\sum_{k=1}^{N}b_{k}q_{k}$ or $n=q_{0}+\sum_{k=1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{1+2t}\neq 0$.
Proof.
From the basic properties of continued fractions we know that $\alpha<1/q_{1}$, and so we have $0<\alpha<2\alpha<\dots<(q_{1}-1)\alpha<q_{1}\alpha<1$.
If $n$ is of one of the two special shapes from the lemma, then we know from Lemma 5.1(2, 8) that $\lVert n\alpha\rVert<\alpha$, and so the intervals of length $\lVert n\alpha\rVert$ (which are either exactly the intervals of length $\left\{n\alpha\right\}$ or their complements) each contain either no points or one point. In particular, the intervals are balanced.
If $n$ is not of one of the two special shapes, then we know from Lemma 5.1(8) that $1/2>\lVert n\alpha\rVert>\alpha$, and so $\left\{n\alpha\right\}>\alpha$.
By simply counting points, we see that the interval $0,\left\{n\alpha\right\})$ contains exactly $\lceil\left\{n\alpha\right\}/\alpha\rceil\geq 2$ points.
On the other hand, since $(q_{1}-1)\alpha<q_{1}\alpha<1$ and $m-1\leq q_{1}-2$, we see that the interval $1-\left\{n\alpha\right\},0)$ contains at most $\lceil\left\{n\alpha\right\}/\alpha\rceil-2$ points. Thus, the intervals are not balanced. ∎
$6.$  Finishing the proof of the full characterisation